OpenStudy (anonymous):

Subtract. 12 – (–17) A. –29 B. –5 C. 5 D. 29

OpenStudy (alexandervonhumboldt2):

12-(-17)=12+17=29

OpenStudy (alexandervonhumboldt2):

-(-)=+

OpenStudy (jareeek):

c?

OpenStudy (anonymous):

-5

OpenStudy (jareeek):

hmmmmm

OpenStudy (alexandervonhumboldt2):

@xXAikaAkihiroXx you are incorrect

OpenStudy (anonymous):

?

OpenStudy (alexandervonhumboldt2):

12-17=-5 but the problem is 12-(-17)

OpenStudy (jareeek):

its positive 5

OpenStudy (alexandervonhumboldt2):

INCOORECT

OpenStudy (anonymous):

A

OpenStudy (jareeek):

aww darn it

OpenStudy (jareeek):

-12 - (-17) Multiply -1 into the parenthesis: -12 + 17 Add: 5

OpenStudy (alexandervonhumboldt2):

wait

OpenStudy (alexandervonhumboldt2):

the problem is 12-(-17) FROM WHERE YOU GOT -12?

OpenStudy (jareeek):

http://brainly.com/question/385281

OpenStudy (jareeek):

you will find the answer

OpenStudy (alexandervonhumboldt2):

you are laugthing on me

OpenStudy (alexandervonhumboldt2):

the problem on brainly is -12-(-17), here the problem is 12-(-17)

OpenStudy (alexandervonhumboldt2):

@AloneS

OpenStudy (anonymous):

D.........gosh yall a bunch of dolts

alones (alones):

yeah @AlexandervonHumboldt2 :) what do need??

OpenStudy (alexandervonhumboldt2):

they seem to laugh on me.......xD

OpenStudy (anonymous):

so what is the answer

OpenStudy (alexandervonhumboldt2):

29

alones (alones):

ha,ha i understand lol @AlexandervonHumboldt2 okay @doopy may i see the question ??

OpenStudy (anonymous):

Subtract. 12 – (–17) A. –29 B. –5 C. 5 D. 29

alones (alones):

yeah @AlexandervonHumboldt2 is right it's D :)

OpenStudy (anonymous):

okay thanks

OpenStudy (jareeek):

do you mind showing me how you got 29

alones (alones):

All the Thank you thingies all go to aleks or what ever his name is :)

alones (alones):

ain't u talking to me @jareeek

OpenStudy (jareeek):

yes

alones (alones):

i'll message u:)

OpenStudy (alexandervonhumboldt2):

@jareeek the problem says `12 – (–17)`. 12-(-17)=12+17=29

OpenStudy (alexandervonhumboldt2):

any more questions?

alones (alones):

Thx @AlexandervonHumboldt2 u took ma question away :) lol well still

OpenStudy (jareeek):

alright i see what happened